HUBBLE SPACE TELESCOPE DAILY REPORT #2170 PERIOD COVERED: 0000Z (UTC) 07/21/98 - 0000Z (UTC) 07/22/98 Daily Status Report as of 203/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/2 7887 (A Survey of Gravitational Lenses as Cosmological Tools II) The Near Infrared Camera (Camera Two) was used to make observations of the gravitational lens candidate system HST16302+8. The propose is to extend our IR survey of gravitational lenses and gravitational lens candidates by obtaining NICMOS images of 14 additional lens systems or candidates. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.5 Completed Three Sets of WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.6 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/CCD/WF/PC-2/NIC/1/2/3 7567 (Supplement to the STIS Dynamics IKP) The Space Telescope Imaging Spectrograph (CCD), the WF/PC-2 and the Near Infrared Camera (Cameras One, Two and Three) were used to make observations of normal galaxy M-87. Detection of supermassive black holes in normal galaxies is one of the grand challenges that HST was designed to meet. A wavelength and CCD flat calibration images were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star HD173739. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.9 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.10 Completed NIC/2/STIS/CCD/WF/PC-2 7220 (Imaging of Quasar Host Galaxies) The Near Infrared Camera (Camera Two), the Space Telescope Imaging Spectrograph (CCD), and the WF/PC-2 were used to make direct imaging observations of the host galaxy for the broad absorption line quasar PG1718+480. The observations were completed as planned, and no anomalies were reported. 1.11 Completed NIC/2 7895 (HST NICMOS SNAPSHOT Observations of Field Galaxies at Z < 1) The Near Infrared Camera (Camera Two) was used to make observations of the field galaxy Gal-141748. The proposed NICMOS observation will provide their morphology in rest frame NIR wavelengths {8000 - 16000 A}, where the effect of dust extinction is less severe, and their near infrared {NIR} colors. The observations were completed as scheduled, and no anomalies were reported. 1.12 Completed NIC/1/2/3 7883 (Infrared Imaging of High Redshift {0.4 <= z <= 1} Tully-Fisher Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the high redshifted Tully- Fisher class galaxy MDS-0305-0. With the advent of 10 meter class telescopes combined with the high spatial resolution of the HST, it is now possible to study the internal kinematics and structural properties of galaxies out to a redshift of z~1. The observations were completed as planned, and no anomalies were reported. 1.13 Completed WF/PC-2 7501 (Expansion Parallax Distances to Planetary Nebulae) The WF/PC-2 was used to make parallax distance measurements of the planetary nebula NGC-7354. The largest impediment to our understanding of planetary nebulae is the lack of reliable distances. Acquired by generalized statistical methods in all but a handful of cases, individual PNe distances are often uncertain by factors of two or more. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was reset five times. The SI C&DH error counter was reset at 202/2130Z per ROP NS-5. The ground system was configured at 202/1927Z to use PRS2 as the prime string. This occurred so that the PRS1/NAS1 port problem recently reported could be investigated. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH